OpenStudy (anonymous):

Find the volume of the region under the graph of f(x,y)=2x+y+1 and above the region y2≤x, 0≤x≤9.

OpenStudy (dumbcow):

I think it would be the following integral \[V = \int\limits_{0}^{9}\int\limits_{-\sqrt{x}}^{\sqrt{x}} (2x+y+1) dydx\]